OpenStudy (nuccioreggie):

Multiply 1 over 3 multiplied by 1 over 5.

OpenStudy (anonymous):

Is this it? \(\dfrac{1}{3} \times \dfrac{1}{5}\)

OpenStudy (7thgrader):

hey igreen i tagged you so whenever ur ready

OpenStudy (anonymous):

@nuccioreggie Is that the question?

OpenStudy (anonymous):

Or \[3^{1}\times5^{1}\]

OpenStudy (nuccioreggie):

@igreen yea

OpenStudy (nuccioreggie):

@igreen. Yea

OpenStudy (anonymous):

Okay, multiply 5 * 3.

OpenStudy (nuccioreggie):

Ok thats 15

OpenStudy (anonymous):

Yes, and what's 1 * 1?

OpenStudy (nuccioreggie):

1

OpenStudy (anonymous):

Yes, so our answer is \(\dfrac{1}{15}\)

OpenStudy (nuccioreggie):

Ok thanks man so can u help me with a couple more please.?

OpenStudy (anonymous):

I would, but I have to go soon..sorry.

OpenStudy (nuccioreggie):

Ok 1more then

OpenStudy (k_lynn):

@nuccioreggie I can help you. Just close the thread and give us a tag.